A collection of dimes and nickels has a total value of $1.65. If there are 20 coins in all, how many of the coins are dimes?
Found 3 solutions by jorel555, Boreal, ikleyn:
Answer by jorel555(1290)

Let d=dimes, then:
.1d+.05(20-d)=1.65
.1d+1-.05d=1.65
.05d=.65
d=.65/.05=13 dimes, 7 nickels!!!!!!!!!!!!!!!
. A solution containing 30% juice is mixed with a solution containing 10% juice to make a 100 gallons of a 12% juice mixture. How much of the juice somution was used
Let n=amount of 30% juice. Then
.3n+.1(100-n)=.12(100)
.3n+10-.1n=12
.2n=2
n=2/.2=10 gallons of 30% juice, 90 gallons of 10% juice!!!!!!!!!!!!!!!

Answer by Boreal(15235)

x=nickel
y=dime
x+y=20
.05x+.10y=1.65
divide first equation by -20
-.05x-0.05y=-1, add them
0.05y=0.65
y=13 dimes
x=7 nickels
That is $1.65

Answer by ikleyn(52778)

.
A collection of dimes and nickels has a total value of $1.65. If there are 20 coins in all, how many of the coins are dimes?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

On detailed solutions of coin problems see the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
in this site.